OpenStudy (anonymous):

answer this question. y=- +b if (4, -10) is on the line

OpenStudy (anonymous):

y=-x+b? or y=-mx+b?

OpenStudy (anonymous):

the question is to figure out what x is, in the blank spot

OpenStudy (anonymous):

okay (x,y) so x=4 and y=-10 plug them in -10=-(4)+b so yeah, you didn't even have to plug it in :)

OpenStudy (anonymous):

thanks:)

OpenStudy (anonymous):

welcome XD